OpenStudy (anonymous):

In triangle ABC, a = 8, b = 6, and C = 60°. Find the measure of A. 16° 46° 74° 106°

OpenStudy (anonymous):

@jim_thompson5910 can you help me please

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

step 1) use the law of cosines to find the length value of c step 2) use law of sines to find angle A

OpenStudy (anonymous):

what would i put for SIN A, SIN B, and SIN C for step 1 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

c^2 = a^2 + b^2 - 2*a*b*cos(C) c^2 = 8^2 + 6^2 - 2*8*6*cos(60) c^2 = 8^2 + 6^2 - 2*8*6*0.5 ... ... ... c = ???

OpenStudy (anonymous):

52?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

c^2 = 52 so c itself is ???

OpenStudy (anonymous):

7.21?? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

we'll leave it as \[\Large \sqrt{52}\]